There are ten pieces of pizza. Anthony ate 1/8 of the pizza for dinner. He ate 1/4 of the pizza for a bedtime snack. How much of the pizza has he eaten in all?
Answer by sofiyac(983)

1/8+1/4=1/8+2/8=3/8 three eighths of a pizza